The market demand for wheat is Q = 100 - 2p + 1pb + 2Y. If the price of wheat, p, is $2, and the price of barley, pb, is $3, and income, Y, is $1000, what is the income elasticity of wheat?


1
Expert's answer
2021-05-24T10:13:54-0400
"Q=100-2\times2+3+2\times1000=1099"

"E=2\times \frac{1000}{1099}=1.81"